OpenStudy (anonymous):

work out 1/5 add 3/7

jhonyy9 (jhonyy9):

1/5 + 3/7 = (7+ 5*3)/35 = (7+15)/35 = 22/35

OpenStudy (anonymous):

1/5 + 3/7 = (7 + 15) / 35 = 22/35

OpenStudy (anonymous):

i got that answer but i was just checking to make sure it was correct and got my 2 makrs in the maths paper :)